OpenStudy (kitten_is_back):

What is the distance between the two points on a number line? Column A Column B 1. X = 4, Y = –17 2. P = 23, Q = 8 3. U = –1, V = –14 4. J = –3, K = 18 A. 13 B. 15 C. 21 Question Resources

OpenStudy (kitten_is_back):

@bullkidd

OpenStudy (kitten_is_back):

@BioHazard9064

OpenStudy (kitten_is_back):

r

OpenStudy (kitten_is_back):

@Vocaloid

OpenStudy (kitten_is_back):

@JoannaBlackwelder

OpenStudy (kitten_is_back):

@BrandonR

OpenStudy (anonymous):

Don't Tag Me In Math Please >.>

OpenStudy (kitten_is_back):

sorry

OpenStudy (kitten_is_back):

@jacobgmiller

OpenStudy (kitten_is_back):

@raiders88

OpenStudy (anonymous):

Let's take 7 , 12 as our first example. We're trying to find out the difference between the two...take 12 minus 7 and you get a difference of 5 between them. Now the first problem, 4, -17...put the larger one first and subtract the smaller one. 4 minus negative 17 is the same as 4 plus 17 which is 21. 23-8 is 15, of course. -1-14 is -15, but the difference between something can't be negative, so it's just 15. 18-(-3) = 18+3 = 21.